OpenStudy (anonymous):

Factor and explain please: 27x^3 + 64y^3

OpenStudy (anonymous):

sum of two cubes \[a^3+b^2=(a+b)(a^2-ab+b^2)\] put \[a=3x, b= 4y\] to get your answer

OpenStudy (anonymous):

(3x)^3 +(4y)^3 =(3x+4y)((3x)^2)-3x.4y+(4y)^2)

OpenStudy (anonymous):

Okay thanks satellite